# Uncertainty-Aware Artificial Intelligence for Gear Fault Diagnosis in Motor Drives thanks: This work was supported by Innovation Fund Denmark through the project AI-Power [17]: Artificial Intelligence for Next-Generation Power Electronics. (Corresponding author: Subham Sahoo)

### v1

@@ V Conclusion
In conclusion, this paper delves into the application of uncertainty-aware AI algorithms for fault diagnosis of gear faults, with a particular focus on Bayesian neural networks (BNNs). Through the utilization of BNNs, we have effectively quantified uncertainties in predictions, providing probabilistic outputs that offer valuable insights into the reliability of diagnostic assessments. We have carried out rigorous test cases by considering different subsets of faults as the training data and identifying reasonable answers from BNN for the unseen faults in a structural way. The effect of noise variance, model parameters and unseen data has been covered in detail with key results based on theoretical foundations. This provides us with a formidable framework for a multi-dimensional decision making process (see Fig. 9) that requires human intervention before finalizing the data-driven algorithm for fault diagnosis of power electronics. This enhances a qualitative assessment of predictions from AI that often goes overlooked due to high accuracy as the sole decision metric. As a follow-up, the decision making process will ascertain corrective actions across the input stage such that further data analysis can be carried out.
The uncertainty estimation and decomposition offer various insights in interpreting the diagnostic results, and issuing an alert for possible wrong diagnosis.
(1) The uncertainty estimation in probabilistic Bayesian deep learning framework is effective in forming the paradigm of trustworthy fault diagnosis. A high level of uncertainty indicates that the model is prone to making incorrect diagnosis and that the results are untrustworthy. The low uncertainty indicates that the input data are the same as or similar to the training data for developing the trained model, the model can be confident in its fault diagnosis. By delivering the uncertainty estimates to human experts, the transparency of these black-box diagnosis models can be improved, and the degree of trustworthiness of model’s diagnosis can be quantified to a greater extent.
(2) Theoretically, epistemic uncertainty comes from the parameters of deep learning models, and represents the deviation of diagnostic results using different weights. Aleatoric uncertainty arises due to the noise in data measurement or unobserved factors, and reflects the inherent stochasticity in the diagnostic results with the fixed model weights.

### v2

@@ V Conclusion
In conclusion, this paper delves into the application of uncertainty-aware AI algorithms for fault diagnosis of gear faults, with a particular focus on Bayesian neural networks (BNNs). Through the utilization of BNNs, we have effectively quantified uncertainties in predictions, providing probabilistic outputs that offer valuable insights into the reliability of diagnostic assessments. We have carried out rigorous test cases by considering different subsets of faults as the training data and identifying reasonable answers from BNN for the unseen faults in a structural way. The effect of noise variance, model parameters and unseen data has been covered in detail with key results based on theoretical foundations. This provides us with a formidable framework for a multi-dimensional decision making process (see Fig. 9) that requires human intervention before finalizing the data-driven algorithm for fault diagnosis of power electronics. This enhances a qualitative assessment of predictions from AI that often goes overlooked due to high accuracy as the sole decision metric. As a follow-up, the decision making process will ascertain corrective actions across the input stage such that further data analysis and investigations can be carried out.
As a result, this work provides a significant leap towards trustworthy machine learning in power electronics by quantifying the uncertainty in the predictions of AI, as a direct measure of either data-driven or model-driven uncertainties.